Need help to simplify square root of 12. I think it is 4 but not sure.
Need help to simplify square root of 75. Not sure at all on this one.

Answer by jim_thompson5910(35256)

sqrt(12) = sqrt(4*3) = sqrt(4)*sqrt(3) = 2*sqrt(3)


So sqrt(12) = 2*sqrt(3)


Use the same idea to simplify the other one.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim